OpenStudy (anonymous):

Consider this system of equations: 23x+35y=12 (equation A) 52y−3x=6 (equation B) The expression that gives the value of x is . The solution for the system of equations is

TheSmartOne (thesmartone):

Welcome to OpenStudy! @helpstudy112 Do you know how to solve by substitution, elimination, or by graphing?

OpenStudy (anonymous):

23x+35y=12;−3x+52y=6 Step: Solve 23x+35y=12 for x: 23x+35y+−35y=12+−35y(Add -35y to both sides) 23x=−35y+12 23x23=−35y+1223(Divide both sides by 23) x=−3523y+1223 Step: Substitute (−3523y+1223) for x in −3x+52y=6: −3x+52y=6 −3(−3523y+1223)+52y=6 130123y+−3623=6(Simplify both sides of the equation) 130123y+−3623+3623=6+3623(Add 36/23 to both sides) 130123y=17423 130123y130123=17423130123(Divide both sides by 1301/23) y=1741301 Step: Substitute (1741301) for y in x=−3523y+1223: x=−3523y+1223 x=−3523(1741301)+1223 x=4141301(Simplify both sides of the equation) Answer: x=414/1301 and y=174/1301

OpenStudy (anonymous):

thank you